Jamie Lee Jackson, age 24, has recently decided to switch from attending college part-time to full-time in order to pursue her business degree, and aims to graduate within the next three years. She has 55 credit hours remaining in order to earn her bachelor's degree, and knows that it will be a challenge to complete her course of study while still working part-time in the bakery department of a local grocery store, where she earns $390 a week. Jamie Lee wants to keep her part-time job at the grocery store as she loves baking and creates very decorative cakes. She dreams of opening her own cupcake café within the next five years.

Jamie Lee currently shares a small apartment with a friend, and they split all of the associated living expenses, such as rent and utilities, although she would really like to eventually have a place of her own. Her car is still going strong, even though it is seven years old, and she has no plans to buy a new one any time soon. She is carrying a balance on her credit card and is making regular monthly payments of $50 with hopes of paying it off within a year. Jamie has also recently taken out a student loan to cover her educational costs and expenses. Jamie Lee just started depositing $1,800 a year in a savings account that earns 2 percent interest, in hopes of having the $9,000 down payment needed to start the cupcake café two years after graduation.

Page 27

Current Financial Situation

Checking account: $1,250

Emergency fund savings account: $3,100

Car: $4,000

Student loan: $5,400

Credit card balance: $400

Gross annual salary: $2,125

Net monthly salary: $1,560

Questions

1.

Using “Your Personal Financial Plan” sheet 2, what are Jamie Lee's short-term financial goals? How do they compare to her intermediate financial goals?

2.

Browse Jamie Lee's current financial situation. Using the SMARTapproach, what recommendations would you make for her to achieve her long-term goals?

3.

Name two opportunity costs that would be considered in Jamie Lee's situation?

4.

Jamie Lee needs to save a total of $9,000 in order to get started in her cupcake café venture. She is presently depositing $1,800 a year in a regular savings account earning 2 percent interest. How much will she have accumulated five years from now in this regular savings account, assuming she will be leaving her emergency fund savings account balance untouched and for a rainy day?
Answer #1
  1. Here and now objectives include: Yearly reserve funds, regularly scheduled installments for her Visa. While the middle objectives include: completing school in three years and sparing $9000 for up front installment of the cupcake bistro. These objectives are connected in that the transient objectives will clear path for the middle of the road objectives.
  2. Jamie Lee's present monetary circumstance is really not all that awful. A few proposals that I would make incorporate the accompanying: Reducing the Emergency finance investment account and increment the yearly funds she is making to begin a cupcake bistro two years after school, begin paying off the understudy advance with the goal that when she begins her business she won't have numerous commitments to meet.
  3. Opportunity costs that Jamie Lee would confront incorporate; losing her week after week procuring of $390 in the event that she stops her low maintenance employment and neglecting to finish school utilizing the 55 credit hours remaining on the off chance that she keeps the activity.
  4. $1800(1.02) = 1836

(1836+1800) (1.02) ^2=3782.89

               (3782.89+1800) (1.02) ^3=5924.61

               (5924.61+1800) (1.02) ^4=8361.37

               (8361.37+1800) (1.02) ^5=11218.97